Zodwa,the operations manager of a large production plant would like to estimate the mean amount of time a worker takes to assemble a new electronic component. Assume that the standard deviation of this assembly time 3.6 minutes.
(a) After observing 120 workers assembling similar devices, Zodwa noticed that their average time was 16.2 minutes. Construct a 92% confidence interval for the mean assembly time.
(b) How many workers should be involved in this study in order to have the mean assembly time estimated up to plus or minus 15 seconds with 92% confidence?

Answer by Boreal(15235)

92% half-interval is z(0.96)*3.6/sqrt(120)=1.75*3.6/sqrt(120)=0.575
the interval is (15.625 min, 16.775 min)
-
error is the half-interval
so e=z(0.96)(sigma/sqrt(n)), error is 0.25 min
0.25=1.75*3.6/sqrt(n)
square both sides and move n over
0.0625n=39.69
n=635.04 or 636 people